Use Lagrange multipliers to find the maximum and minimum values of f(x,y)=x−2y subject to the constraint x^2+y^2=5, if such values exist.

maximum = ?

minimum = ?
1
Expert's answer
2020-04-08T12:28:59-0400

Given "f(x,y) =x-2y"

Let  "g(x,y) = x^{2}+y^{2}-5" and

let "F = f(x,y)+\lambda g(x,y)"

where "\lambda"  is the Lagrangian multiplier.

"F = x - 2y + \lambda(x^{2} + y^{2} -5)"


To find the maximum and minimum values, we have to solve the system

"\dfrac{\partial F}{\partial x} = 0"

"\dfrac{\partial F}{\partial y} = 0" and 

"\dfrac{\partial F}{\partial \lambda} = 0"


We get,

"1 + 2x\lambda = 0\\ -2 + 2y\lambda = 0\\ x^{2} + y^{2} = 5."

From the first and the second equation we get,

"x = -\dfrac{1}{2\lambda}\\ y = \dfrac{1}{\lambda}\\"

Using these values in the third equation we get,

"\dfrac{1}{4\lambda^{2}} + \dfrac{1}{\lambda^{2}} = 5\\ \lambda^{2} = \dfrac{1}{4}\\ \lambda = \pm\dfrac{1}{2}\\"


When "\lambda = -\dfrac{1}{2}",

"x= 1, y =-2"

When "\lambda = \dfrac{1}{2}",

"x= -1, y =2"

The maximum and minimum values of f(x,y) are respectively,

"f(1,-2) = 1+4 = 5\\ f(-1,2) = -1-4=-5"